# Lucas Resende CREST, ENSAE, Institut Polytechnique de Paris, France Guillaume Lecué ESSEC Business School, France Lionel Wilner CREST, ENSAE, Institut Polytechnique de Paris, France Philippe Choné CREST, ENSAE, Institut Polytechnique de Paris, France

### v1

@@ 4.1 Consistency
Assumption 2 is not restrictive: as $n^{(k)}$ grows, the number of positive entries $Y_{\bf i}^{(k)}$ with disjoint indices should also increase, and therefore, the number of pairs of active polyads sharing no edge should outnumber those sharing at least one edge. Notice that if one assumes a distribution on the fixed effects and on the covariates this assumption is automatically true, as in this case the probability that $\xi$ and $\xi^{\prime}$ are both active is always upper and lower bounded by a constant, thus it is just a matter of comparing the number of pairs of polyads – which has order $(n^{(k)})^{4}$ – and the number of pairs of polyads sharing an edge – which has order $(n^{(k)})^{3}$. This last approach is used by graham2017econometric and jochmans2018semiparametric.
@@
The proof can be found in Section C.3. The proof uses the convexity of the loss function as a key ingredient. It is based on a slight modification of Theorem 2.7 of newey1994large that requires revisiting classical results in convex analysis (such as those from Chapter 10 in rockafellar-1970a), as well as some results in asymptotic statistics from newey1994large and Andersen1982. ∎
@@ 4.2 Asymptotic normality
The last approximation shows that to control the fluctuations of $\widehat{\beta}^{(k)}-\beta_{\star}$ we need to control the fluctuations of $\nabla\widehat{Q}^{(k)}\left(\beta_{\star}\right)$, which is not a sum of independent random variables. Following graham2017econometric; jochmans2018semiparametric, we use Hájek projections to approximate $\nabla\widehat{Q}^{(k)}\left(\beta_{\star}\right)$ by a sum of independent random variables for which classical central limit theorems hold. The following assumption is useful to control the error of this approximation:
@@
The proof of Theorem 2, given in Section B.2, follows from a general lemma presented in Section B.1, adapted from Brunel’s lecture notes victor, which reference MR1026303 and MR1186263. Unlike the classical setup, where the loss is a sum of independent terms, we handle dependent polyads and do not assume the existence of a limiting $Q_{\infty}$, only a limiting covariance matrix at $\beta_{\star}$. Hence, we extend the standard asymptotic normality proof under convexity to this dependent setting. ∎
@@ 5 Computational implementation
Lemma 1 provides the tools to solve the optimization problem (14) defining the polyads estimator $\widehat{\beta}_{\Xi}$. Since each term in the loss function is convex and we have expressions for both their gradients and their Hessians, we can efficiently solve the minimization problem using Newton’s method. However, computing the gradient and Hessian involves summing over all polyads $\xi\in\Xi$, which naively requires looping over $\prod_{d=1}^{D}n_{d}(n_{d}-1)$ terms, an operation that quickly becomes computationally expensive. The main goal of this section is to reduce this complexity by avoiding unnecessary iterations, which is done characterizing the set of active polyads $\Xi_{a}$. Besides that, we also discuss the implementation of two approximations of the variance. These approximations are essential to construct confidence intervals. The methods presented in this section are of special interest when $Y$ is sparse. In particular, the computational complexity of our methods outperforms the PPML alternative correia2019ppmlhdfe when the size of $E=\{{\bf i}:Y_{\bf i}>0\}$ is of order smaller than $\sqrt{n}$.
@@ 5.5 Final cost analysis and practical considerations
In practice, Newton’s method converges in less than $10$ iterations, yielding $O(|E|^{2})$ operations. This quantity is to be compared with the fast implementation of PPML from correia2019ppmlhdfe, which is $O(n)$. Our analysis suggests that our method is faster when $|E|\ll\sqrt{n}$ and competitive when $|E|$ is of order $\sqrt{n}$. Although being the standard practice when reporting the complexity of algorithms, the big-O notation hides a constant that matters to practitioners. In the next section, we empirically verify that, as predicted by our cost analysis, our method outperforms PPML in terms of computational time when $|E|$ is smaller than $\sqrt{n}$ and remains competitive as $|E|$ grows. Indeed, in our computational setup (see Section 6) the running time of our method is shorter than that of PPML as long as $|E|\leq 15\sqrt{n}$. This is, for example, the case of a bipartite network with $n_{1}=n_{2}$ such that the average degree of a node is at most $15$.
@@
We also note that the improvements developed in this section, and particularly the construction of $\Xi_{a}^{\star}$, may be useful for other polyads-based methods. For example, jochmans2017two proposes a generalized method of moments for two-way models with $n_{1}=n_{2}=\sqrt{n}$. His tetrad-based estimator leverages matrix-multiplication tricks and achieves a computational complexity of order $O(n^{1.1877})$ when using the best available theoretical matrix-multiplication algorithm. In contrast, our $O(|E|^{2})$ complexity can yield substantial gains in sparse regimes.
@@ 6 Experiments
We provide experiments comparing our method with PPML and with the analytical debias proposed by zylkin2024bootstrap. We consider artificial and real data. With artificial data we investigate the impact of the incidental parameter problem while knowing the correct value of $\beta_{\star}$. With real data we display evidence of the incidental parameter bias and show how it may lead towards wrong conclusions in inference. We also discuss the computational time and the effect of sparsity for all methods.
@@ 6.1 Artificial data
We conduct computational experiments using a three-way data-generating process inspired by weidner2021bias; see Example 3. The dimensions are $n_{1}=n_{2}$ (varied) and $n_{3}=5$ (fixed). The fixed effects $u_{ij},w_{it},v_{jt}$ are i.i.d. $\mathcal{N}(0,1/16)$, and $\beta_{\star}=1$. The covariates $X_{ijt}$ are correlated with the fixed effects and, along the third axis, with their own past values:
@@
We compare three estimators: PPML, PPML debiased, and our polyads estimator. For PPML we use the fast implementation of correia2019ppmlhdfe; for PPML debiased we use the analytical correction of weidner2021bias via their Stata package. Our first experiment varies the graph density with
@@ 6.2 Real data
Given the strong policy concern about care accessibility, it is important to understand how the reform has affected potential patients. Have certain categories of patients (patients with chronic diseases, women, elderly patients, patients living in medically undeserved areas, low-income patients, etc.) been disproportionately affected by the reform? Has the reform caused patients to seek care further away from their home? To answer these questions, we need to perform the analysis at the doctor-patient level rather than at the doctor level. Using a three-way model and controlling for dyads fixed effects allows to estimate how the reform has changed the doctor-patient connections. To illustrate the method, we highlight in this subsection the role of two variables: patient sex and geographic distance (spatial accessibility).
The outcome $Y_{ijt}$ is the number of visits by patient $i$ to doctor $j$ on month $t$. We aggregate data at the city-sex level. The high number of potential patients in each city-sex group makes to the Poisson assumption plausible. [^9] The aggregate number of consultations in each group is the sum of (possibly heterogeneous) Bernoulli distributions that represent the occurrence of a consultation for all potential patients in the group. This sum follows approximately a Poisson distribution under the conditions exposed in le1960approximation. The assumption that the individual occurrences of a consultation are independent across potential patients is relaxed in galambos1973general and serfling1978some. The index $i$ (resp. $j$) stands thus for the set of patients (resp. doctors) in a given municipality with same gender. The treatment $T_{j}$ is a binary variable equal to 1 for sector 1 GPs, and to 0 for direct access specialists. The reform has been implemented from May 2017 onward, hence the definition of $\text{Post}_{t}$, a dummy variable equal to 1 after that date. The model includes three features that account for any possible policy-induced change in homophily preferences as regards gender and spatial dimensions: the interactions between $\text{Post}_{t}\times T_{j}$ (as in any difference-in-differences approach) and (i) a dummy variable equals to 1 when patients and doctors have the same sex, (ii) a dummy variable that is equal to 1 when the medical office chosen is located in the same municipality as the patient’s home, and (iii) travel time (measured in minutes between the centroids of municipalities).
@@
We compare three estimators: PPML, PPML debiased and our polyads estimator. Because the full dataset contains $n_{1}=69{,}265$ patient groups, $n_{2}=16{,}941$ doctors, and $n_{3}=34$ months — corresponding to roughly $n\approx 40$ billion edges and |E|=56,034,015|E|=56{,}034{,}015 — direct estimation on the full graph is computationally infeasible for all methods. We therefore adopt a subsampling strategy combined with meta-analysis.
For each configuration, we sample a proportion $s\in\{2\%,3\%,4\%\}$ of patient groups and the same proportion of doctors, while always retaining all $34$ months [^10] Notice that we sample the groups of patients, the doctors and then get all edges through the sampled nodes, keeping all times. This sampling strategy is consistent with Assumption 1 and also with the sampling assumptions in fernandez2016individual; weidner2021bias; graham2017econometric; jochmans2018semiparametric. One could also propose sampling directly the edges and not the nodes, but there is evidence that this procedure can lead to bias, see subsamplegraphs for details.. Each subsample consists of independent random draws of patient and doctor groups; doctors in the treatment and control groups are sampled independently to ensure comparability. We repeat the procedure independently to obtain $100$ subsamples. The final estimates are obtained via a random-effects meta-analysis using the default implementation in the statsmodels Python library, which is based on the iterated method of paule1982consensus, a refinement of the classical method of dersimonian1986meta.
@@
Table 1: Random effects meta-analysis estimates with 95% confidence intervals (in parentheses) by subsample proportion and method. $\beta_{\texttt{d}}$ estimates are scaled by $10^{4}$; $\beta_{\texttt{sc}}$ and $\beta_{\texttt{ss}}$ by $10^{2}$.
@@
The results are the following. First, Table 1 and Figure 5 show that debiased PPML has many outliers, mostly on ‘small’ samples (admittedly, not that small since $|E|$ amounts to about 20,000 with the 2% subsample size), which results in (very) large confidence intervals and point estimates. As expected, the variance decreases with the subsample size, regardless of the estimation method used.
Second, the running time of PPML and debiased PPML scales with $n$ while the polyads time scales with $|E|^{2}$ (Table 2).
@@ Appendix A Consistency of $\widehat{\beta}_{\Xi}$ in the Poisson model
The proof of the consistency result uses the convexity of the loss function as a key ingredient. It is based on a slight modification of Theorem 2.7 of newey1994large that requires to revisit some classical results in convex analysis (such as those from Chapter 10 in rockafellar-1970a) as well as some results in asymptotic statistics from newey1994large and Andersen1982. All these revisited versions of these classical results may be found in Section C.3. Theorem 1 is a consequence of Lemma A.1 below whose proof may be found also in Section C.3.
@@
They are two main advantages of Lemma A.1: first, as in Theorem 2.7 from newey1994large, we don’t need to assume that $\Theta$ is compact (that will be useful for us since we are minimizing over all $\beta\in\mathbb{R}^{p}$ to define the polyads estimator); second, we don’t need to assume that the sequence of risk functions $(Q_{N})_{N}$ has a limit or that it is the same function for all $N$. The latter assumption is problematic when the fixed effects are considered as parameters (or when they are random variables and we work conditionally to them); the same remark holds for the co-variable vectors $(X_{{\bf i}})_{\bf i}$. This is our main motivation to use Lemma A.1 in place of the classical Theorem 2.7 from newey1994large where a limiting function is assumed to exists.
@@ B.1 Main theorem for asymptotic normality under convexity assumption
The following result is adapted from Victor-Emmanuel Brunel lecture notes victor who cites MR1026303 and MR1186263 as classical references for asymptotic statistics under convexity. However, all the later results use a loss function which is a sum of independent variables. This is not our case here since our loss function sums over all polyads that are not necessarily independent, moreover, as mentioned previously in the section on the consistency result above, we do not assume the existence of a limiting risk function; we only require the existence of a limiting covariance matrix of the risk in $\beta_{\star}$. We therefore need to adapt the classical proof of asymptotic normality under convexity assumption to our setup. The proof of asymptotic normality of the polyads estimator will rely on the following general result.
@@
The latter result holds pointwise (i.e. for a given $t$) but we know that a sequence of convex functions that converges pointwise also converges uniformly over all compact set (see Chapter 10 in rockafellar-1970a). However, we cannot use directly this result in our case since we are in a situation where it is the difference of two sequences of convex functions that converges pointwise to zero in (B.1). Hence, as in the proof of consistency of the polyads estimator from the previous section, we apply the convergence result from Lemma C.5. As a consequence, it follows from the pointwise convergence (B.1) that on the compact set $B_{2}(0,2\varepsilon_{0})$, we have

### v2

@@ 4.1 Consistency
Assumption 2 is not restrictive: as $n^{(k)}$ grows, the number of positive entries $Y_{\bf i}^{(k)}$ with disjoint indices should also increase, and therefore, the number of pairs of active polyads sharing no edge should outnumber those sharing at least one edge. Notice that if one assumes a distribution on the fixed effects and on the covariates this assumption is automatically true, as in this case the probability that $\xi$ and $\xi^{\prime}$ are both active is always upper and lower bounded by a constant, thus it is just a matter of comparing the number of pairs of polyads – which has order $(n^{(k)})^{4}$ – and the number of pairs of polyads sharing an edge – which has order $(n^{(k)})^{3}$. This last approach is used by Graham (2017) and Jochmans (2018).
@@
The proof can be found in Section C.3. The proof uses the convexity of the loss function as a key ingredient. It is based on a slight modification of Theorem 2.7 of Newey and McFadden (1994) that requires revisiting classical results in convex analysis (such as those from Chapter 10 in Rockafellar (1970)), as well as some results in asymptotic statistics from Newey and McFadden (1994) and Andersen and Gill (1982). ∎
@@ 4.2 Asymptotic normality
The last approximation shows that to control the fluctuations of $\widehat{\beta}^{(k)}-\beta_{\star}$ we need to control the fluctuations of $\nabla\widehat{Q}^{(k)}\left(\beta_{\star}\right)$, which is not a sum of independent random variables. Following Graham (2017); Jochmans (2018), we use Hájek projections to approximate $\nabla\widehat{Q}^{(k)}\left(\beta_{\star}\right)$ by a sum of independent random variables for which classical central limit theorems hold. The following assumption is useful to control the error of this approximation:
@@
The proof of Theorem 2, given in Section B.2, follows from a general lemma presented in Section B.1, adapted from Brunel’s lecture notes Brunel (2020), which reference Haberman (1989) and Niemiro (1992). Unlike the classical setup, where the loss is a sum of independent terms, we handle dependent polyads and do not assume the existence of a limiting $Q_{\infty}$, only a limiting covariance matrix at $\beta_{\star}$. Hence, we extend the standard asymptotic normality proof under convexity to this dependent setting. ∎
@@ 5 Computational implementation
Lemma 1 provides the tools to solve the optimization problem (14) defining the polyads estimator $\widehat{\beta}_{\Xi}$. Since each term in the loss function is convex and we have expressions for both their gradients and their Hessians, we can efficiently solve the minimization problem using Newton’s method. However, computing the gradient and Hessian involves summing over all polyads $\xi\in\Xi$, which naively requires looping over $\prod_{d=1}^{D}n_{d}(n_{d}-1)$ terms, an operation that quickly becomes computationally expensive. The main goal of this section is to reduce this complexity by avoiding unnecessary iterations, which is done characterizing the set of active polyads $\Xi_{a}$. Besides that, we also discuss the implementation of two approximations of the variance. These approximations are essential to construct confidence intervals. The methods presented in this section are of special interest when $Y$ is sparse. In particular, the computational complexity of our methods outperforms the PPML alternative Correia et al. (2019) when the size of $E=\{{\bf i}:Y_{\bf i}>0\}$ is of order smaller than $\sqrt{n}$.
@@ 5.5 Final cost analysis and practical considerations
In practice, Newton’s method converges in less than $10$ iterations, yielding $O(|E|^{2})$ operations. This quantity is to be compared with the fast implementation of PPML from Correia et al. (2019), which is $O(n)$. Our analysis suggests that our method is faster when $|E|\ll\sqrt{n}$ and competitive when $|E|$ is of order $\sqrt{n}$. Although being the standard practice when reporting the complexity of algorithms, the big-O notation hides a constant that matters to practitioners. In the next section, we empirically verify that, as predicted by our cost analysis, our method outperforms PPML in terms of computational time when $|E|$ is smaller than $\sqrt{n}$ and remains competitive as $|E|$ grows. Indeed, in our computational setup (see Section 6) the running time of our method is shorter than that of PPML as long as $|E|\leq 15\sqrt{n}$. This is, for example, the case of a bipartite network with $n_{1}=n_{2}$ such that the average degree of a node is at most $15$.
@@
We also note that the improvements developed in this section, and particularly the construction of $\Xi_{a}^{\star}$, may be useful for other polyads-based methods. For example, Jochmans (2017) proposes a generalized method of moments for two-way models with $n_{1}=n_{2}=\sqrt{n}$. His tetrad-based estimator leverages matrix-multiplication tricks and achieves a computational complexity of order $O(n^{1.1877})$ when using the best available theoretical matrix-multiplication algorithm. In contrast, our $O(|E|^{2})$ complexity can yield substantial gains in sparse regimes.
@@ 6 Experiments
We provide experiments comparing our method with PPML and with the analytical debias proposed by Zylkin (2024). We consider artificial and real data. With artificial data we investigate the impact of the incidental parameter problem while knowing the correct value of $\beta_{\star}$. With real data we display evidence of the incidental parameter bias and show how it may lead towards wrong conclusions in inference. We also discuss the computational time and the effect of sparsity for all methods.
@@ 6.1 Artificial data
We conduct computational experiments using a three-way data-generating process inspired by Weidner and Zylkin (2021); see Example 3. The dimensions are $n_{1}=n_{2}$ (varied) and $n_{3}=5$ (fixed). The fixed effects $u_{ij},w_{it},v_{jt}$ are i.i.d. $\mathcal{N}(0,1/16)$, and $\beta_{\star}=1$. The covariates $X_{ijt}$ are correlated with the fixed effects and, along the third axis, with their own past values:
@@
We compare three estimators: PPML, PPML debiased, and our polyads estimator. For PPML we use the fast implementation of Correia et al. (2019); for PPML debiased we use the analytical correction of Weidner and Zylkin (2021) via their Stata package. Our first experiment varies the graph density with
@@ 6.2 Real data
Given the strong policy concern about care accessibility, it is important to understand the impact of the reform on potential patients. Have certain categories of patients (patients with chronic diseases, women, elderly patients, patients living in medically undeserved areas, low-income patients, etc.) been disproportionately affected by the reform? Has the reform caused patients to seek care further away from their home? To answer these questions, we need to perform the analysis at the doctor-patient level rather than at the doctor level. Using a three-way model and controlling for dyads fixed effects allows to estimate how the reform has changed the doctor-patient connections. To illustrate the method, we highlight in this subsection the role of two variables: gender and geographic distance (spatial accessibility).
The outcome $Y_{ijt}$ is the number of visits by patient $i$ to doctor $j$ on month $t$. We aggregate data at the city-sex level. The high number of potential patients in each city-sex group makes to the Poisson assumption plausible. [^9] The aggregate number of consultations in each group is the sum of (possibly heterogeneous) Bernoulli distributions that represent the occurrence of a consultation for all potential patients in the group. This sum follows approximately a Poisson distribution under the conditions exposed in Le Cam (1960). The assumption that the individual occurrences of a consultation are independent across potential patients is relaxed in Galambos (1973) and Serfling (1978). The index $i$ (resp. $j$) stands thus for the set of patients (resp. doctors) in a given municipality with given gender. The treatment $T_{j}$ is a binary variable equal to 1 for sector 1 GPs, and to 0 for direct access specialists. The reform has been implemented from May 2017 onward, hence the definition of $\text{Post}_{t}$, a dummy variable equal to 1 after that date. The model includes three features that account for any possible policy-induced change in homophily preferences as regards gender and spatial dimensions: the interactions between $\text{Post}_{t}\times T_{j}$ (as in any difference-in-differences approach) and (i) a dummy variable equals to 1 when patients and doctors have the same sex, (ii) a dummy variable that is equal to 1 when the medical office chosen is located in the same municipality as the patient’s home, and (iii) travel time (measured in minutes between the centroids of municipalities).
@@
We compare three estimators: the standard PPML estimator, the analytical correction of Weidner and Zylkin (2021), and our polyads estimator. Because the full dataset contains $n_{1}=69{,}265$ patient groups, $n_{2}=16{,}941$ doctors, and $n_{3}=34$ months — corresponding to roughly $n\approx 40$ billion edges and |E|=56,034,015|E|=56{,}034{,}015 — direct estimation on the full graph is computationally infeasible for all methods. We therefore adopt a subsampling strategy combined with meta-analysis.
For each configuration, we sample a proportion $s\in\{2\%,3\%,4\%\}$ of patient groups and the same proportion of doctors, while always retaining all $34$ months [^10] Notice that we sample the groups of patients, the doctors and then get all edges through the sampled nodes, keeping all times. This sampling strategy is consistent with Assumption 1 and also with the sampling assumptions in Fernández-Val and Weidner (2016); Weidner and Zylkin (2021); Graham (2017); Jochmans (2018). One could also propose sampling directly the edges and not the nodes, but there is evidence that this procedure can lead to bias, see Shalizi and Rinaldo (2013) for details.. Each subsample consists of independent random draws of patient and doctor groups; doctors in the treatment and control groups are sampled independently to ensure comparability. We repeat the procedure independently to obtain $100$ subsamples. The final estimates are obtained via a random-effects meta-analysis using the default implementation in the statsmodels Python library, which is based on the iterated method of Paule and Mandel (1982), a refinement of the classical method of DerSimonian and Laird (1986).
@@
Table 1: Random effects meta-analysis estimates with 95% confidence intervals (in parentheses) by subsample proportion and method (100 replications). $\beta_{\texttt{d}}$ estimates are scaled by $10^{4}$; $\beta_{\texttt{sc}}$ and $\beta_{\texttt{ss}}$ by $10^{2}$.
@@
The results are the following. First, Table 1 and Figure 5 show that debiased PPML has many outliers, mostly on ‘small’ samples (admittedly, not that small since $|E|\approx 20,000$ with the 2% sampling), which results in (very) large confidence intervals and point estimates. As expected, the variance decreases with the subsample size, regardless of the estimation method used.
Second, as reported on Table 2, the running time of PPML and debiased PPML scales with $n$ while the polyads time scales with $|E|^{2}$.
@@
Code resources
A fast python-based implementation of the polyads method here presented is available at the Github repository lucasresenderc/polyads, available at https://github.com/lucasresenderc/polyads.
Acknowledgments
We are grateful to Clément de Chaisemartin, Laurent Davezies, Xavier D’Haultfœuille, Francis Kramarz, and Yannick Guyonvarch for helpful comments. We thank the Agence Nationale de la Recherche for financial support (ANR-23-CE36-0014).
@@ Appendix A Consistency of $\widehat{\beta}_{\Xi}$ in the Poisson model
The proof of the consistency result uses the convexity of the loss function as a key ingredient. It is based on a slight modification of Theorem 2.7 of Newey and McFadden (1994) that requires to revisit some classical results in convex analysis (such as those from Chapter 10 in Rockafellar (1970)) as well as some results in asymptotic statistics from Newey and McFadden (1994) and Andersen and Gill (1982). All these revisited versions of these classical results may be found in Section C.3. Theorem 1 is a consequence of Lemma A.1 below whose proof may be found also in Section C.3.
@@
They are two main advantages of Lemma A.1: first, as in Theorem 2.7 from Newey and McFadden (1994), we don’t need to assume that $\Theta$ is compact (that will be useful for us since we are minimizing over all $\beta\in\mathbb{R}^{p}$ to define the polyads estimator); second, we don’t need to assume that the sequence of risk functions $(Q_{N})_{N}$ has a limit or that it is the same function for all $N$. The latter assumption is problematic when the fixed effects are considered as parameters (or when they are random variables and we work conditionally to them); the same remark holds for the co-variable vectors $(X_{{\bf i}})_{\bf i}$. This is our main motivation to use Lemma A.1 in place of the classical Theorem 2.7 from Newey and McFadden (1994) where a limiting function is assumed to exists.
@@ B.1 Main theorem for asymptotic normality under convexity assumption
The following result is adapted from Victor-Emmanuel Brunel lecture notes Brunel (2020) who cites Haberman (1989) and Niemiro (1992) as classical references for asymptotic statistics under convexity. However, all the later results use a loss function which is a sum of independent variables. This is not our case here since our loss function sums over all polyads that are not necessarily independent, moreover, as mentioned previously in the section on the consistency result above, we do not assume the existence of a limiting risk function; we only require the existence of a limiting covariance matrix of the risk in $\beta_{\star}$. We therefore need to adapt the classical proof of asymptotic normality under convexity assumption to our setup. The proof of asymptotic normality of the polyads estimator will rely on the following general result.
@@
The latter result holds pointwise (i.e. for a given $t$) but we know that a sequence of convex functions that converges pointwise also converges uniformly over all compact set (see Chapter 10 in Rockafellar (1970)). However, we cannot use directly this result in our case since we are in a situation where it is the difference of two sequences of convex functions that converges pointwise to zero in (B.1). Hence, as in the proof of consistency of the polyads estimator from the previous section, we apply the convergence result from Lemma C.5. As a consequence, it follows from the pointwise convergence (B.1) that on the compact set $B_{2}(0,2\varepsilon_{0})$, we have